Let's look at another one of homer's rocket launches. It was launched from ground level with an initial velocity of 208 feet per second. its distance in feet from the ground after t seconds is given by S(t)=-16t^2+208t.
What is the maximum altitude (height) the rocket ail attain during its flight?

Answer by stanbon(75887)

Let's look at another one of homer's rocket launches. It was launched from ground level with an initial velocity of 208 feet per second. its distance in feet from the ground after t seconds is given by S(t)= -16t^2+208t.
What is the maximum altitude (height) the rocket ail attain during its flight?
------
Max occurs when t = -b/(2a) = -208/(-32) = 6.5 seconds
----
Max height is S(6.5) = 676 ft.
---------------
Cheers,
Stan H.
-------------